When a horizontal of 30 N acts upon a 5 kg object on a surface the object moves with constant velocity. Find the coefficient of kinetic friction, between the object and the surface (gravity is equal to 10)
1
Expert's answer
2013-02-19T09:43:07-0500
The object moves with constant velocity, therefore its acceleration equals 0:
X-axis: F - F_fr = ma = 0 => F = F_fr.
F - a horizontal force& (30 N)
F_fr - friction force
Y-axis: N = mg &
F_fr = k*N & => F_fr = k*mg
k - the coefficient of kinetic friction
Therefore:
F = k*mg => k=F/(mg)
k=F/(mg) = 30/(5*10) = 3/5 = 0.6
Answer: the coefficient of kinetic friction equals 0.6